OpenStudy (anonymous):

find the length of the curve y=x^(3/2) from x=0 to x=4

OpenStudy (anonymous):

please someone help! I got 512/729 is that right?

terenzreignz (terenzreignz):

How?

OpenStudy (anonymous):

lol i did it over and got 8...

OpenStudy (anonymous):

\[\int\limits_{0}^{4}\sqrt{1+\frac{ 9x }{ 4 }}\]

OpenStudy (anonymous):

thats the equation. i know i got that part right its the algebra i keep making mistakes on

terenzreignz (terenzreignz):

You set up the definite integral correctly, but there's a bit of an error in how you evaluated it...

terenzreignz (terenzreignz):

\[\int\limits_{0}^4 \sqrt{1+\frac{9x}4} \ dx\] So, I take it you took \[\Large u = 1+\frac{9x}4\]\[\Large du = \frac94 dx\]

OpenStudy (anonymous):

yeah. after integrating i got \[\frac{ 8 }{ 27 } ( 1 + \frac{ 27x ^{\frac{ 3 }{ 2 }} }{ 8 }) \]

OpenStudy (anonymous):

after evaluating everything in u to the 3/2 power.

terenzreignz (terenzreignz):

\[\frac49\int\limits_{1}^{10} \sqrt{u} \ du\]

OpenStudy (anonymous):

whoah how did the inegral change from 0-4 to 1-10?

terenzreignz (terenzreignz):

Because that is u evaluated at the points x = 0 and x = 4

OpenStudy (anonymous):

can you please just do it out completely so i can see how you're getting that?

terenzreignz (terenzreignz):

I just replaced the 1+ (9x/4) with u and the dx with the (4/9)du

OpenStudy (anonymous):

what do you get as an answer?

OpenStudy (anonymous):

whenever i do it i either get 8 or 512/729. the last one looks off cause that means it isnt even 1 unit long but 8 seems too big